OpenStudy (anonymous):

PLEASE HELP!!!! How to Prove: Sin A - 2Sin^3 A/2Cos^3 A - Cos A = Tan A

OpenStudy (bahrom7893):

Wait what did you mean? (Sin A - 2Sin^3 A) / (2Cos^3 A - Cos A) or Sin A - (2Sin^3 A/2Cos^3 A) - Cos A? Please use parenthesis..

OpenStudy (anonymous):

i guess the first one is correct, dont know it is equal to TanA or not

OpenStudy (bahrom7893):

Let me see: (Sin A - 2Sin^3 A) / (2Cos^3 A - Cos A)= Sin A [1 - 2Sin^2 A] / Cos A [ 2Cos^2 A - 1]

OpenStudy (anonymous):

Yeah it is (Sin A - 2Sin^3 A) / (2Cos^3 A - Cos A) = Tan A

OpenStudy (bahrom7893):

Tan A = Sin A / Cos A, so now we have to prove that the rest is equal to one. So: [1 - 2Sin^2 A] / [ 2Cos^2 A - 1]

OpenStudy (bahrom7893):

working this out on paper..

OpenStudy (bahrom7893):

[1 - 2Sin^2 A] / [ 2Cos^2 A - 1] = [1 - 2 ( 1 - Cos^2 A )] / [ 2Cos^2 A - 1]

OpenStudy (bahrom7893):

[1 - 2 ( 1 - Cos^2 A )] / [ 2Cos^2 A - 1] = [1 - 2 + 2Cos^2 A )] / [ 2Cos^2 A - 1] = [-1 + 2Cos^2 A )] / [ 2Cos^2 A - 1] =

OpenStudy (bahrom7893):

[ 2Cos^2 A - 1] / [ 2Cos^2 A - 1] = 1 So there you go, it turns out to be Tan A * 1

OpenStudy (bahrom7893):

Do I get good answer? haha